OpenStudy (anonymous):

will medal and fan. 30 is 12% of what number? ? ----

OpenStudy (anonymous):

to solve the question replace ( 12/100 ) * X = 30 X is the number you look for 12% = 12/100

OpenStudy (anonymous):

so 12/100?

OpenStudy (anonymous):

is above of so : 30/x=12/100 3000=12x divide by 12 x=250

OpenStudy (anonymous):

oh wait i did it backwards

OpenStudy (anonymous):

yes now medal

OpenStudy (anonymous):

I am sorry @liljohn10. I do not believe that 12/100 is correct

OpenStudy (anonymous):

I was checking my answer on here because I thought it was 250 but I wanted to make sure I was correct but it seems that @l_o_w_key answer is more practical. 250

OpenStudy (anonymous):

wait one sec

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

yeah I'm right

OpenStudy (anonymous):

OK thank you I already gave you your medal

OpenStudy (anonymous):

thank you!